MICROWAVE INVESTIGATION OF THE GROUND AND FIRST EXCITED TORSIONAL STATES OF ACETIC ACID MONOHYDRATE.

\underline{GALEN SEDO}, JAMIE DORAN, SHENGHAI WU, KENNETH R. LEOPOLD, University of Minnesota, Department of Chemistry, 207 Pleasant St SE, Minneapolis MN 55455.

Rotational spectra of CH_{3}COOH-H_{2}O and ^{13}CH_{3}COOH-H_{2}O have been observed by Fourier transform microwave spectroscopy. A total of 14 A state and 13 E state transitions have been analyzed. The resulting rotational constants are consistent with a structure similar to that of acetic acid dimer, where the water moiety forms a primary hydrogen bond with the carboxylic acid proton and a secondary hydrogen bond with the carbonyl oxygen. The barrier to internal rotation of the methyl group was determined to be 138.45(3) cm^{−1}, which represents a decrease of 18.7% when compared with that of the acetic acid monomer.